Hardware move

The hardware move has gone relatively smoothly today. We’ve had some configuration issues that prevented some Z39.50 users from pulling back records and another configuration problem that meant a small percentage of the records weren’t visible. That should all be fixed now, but if you see something else that looks like a problem, then please let us know.

The DNS entry for copac.ac.uk was changed at about 10am this morning. At 4pm we’re still seeing some usage on the old hardware. However, most usage started coming through to the new machine very soon after the DNS change.

The change over to the new hardware has involved a lot of preparation over many weeks. Now it’s done we can now get back to re-engineering Copac… a new database backend and new search facilities for the users.